# EPN: An Ego Vehicle Planning-Informed Network for Target Trajectory Prediction

### v1

@@ III-C LSTM Decoder
In the LSTM decoder module, the endpoint encoder is used to encode the corrected endpoint $\widehat{G}$. The resulting encoded feature vector is then concatenated with the environmental feature vector and input into the LSTM decoder. This process generates complete trajectories for each corrected endpoint $\widehat{G}$, resulting in multi-modal trajectory prediction outcomes, the calculation process is shown in formulas (25) and (26):
$refined\_end\_feature={E_{end}}(\widehat{G})$ (25)
$\widehat{Y}=LSTM(concat(enc,refined\_end\_feature))$ (26)
@@
Due to the significant variations in the absolute coordinate values of vehicles across different scenarios, we aim to mitigate the adverse effects of these absolute values on the performance of trajectory prediction models. To achieve this, we predict the relative displacement $\left({\delta x_{t+\tau}^{i},\delta y_{t+\tau}^{i}}\right)$ of the target vehicle over the next ${T_{f}}$ time steps, where $\tau\in\left({1,2,...,{T_{f}}}\right)$. By predicting the relative displacement, we can improve the accuracy of the predictions. Once the relative displacement is predicted, the absolute coordinates of the predicted trajectory can be calculated using the formulas (27) and (28):
@@
III-D Loss Function
We use the trajectory prediction error ${L_{pred}}$ and the conditional variational autoencoder error ${L_{cvae}}$ as loss functions. The trajectory prediction error ${L_{pred}}$ is calculated as the mean squared error (MSE) between the predicted trajectory $\widehat{Y}$ and the true trajectory $Y$, as well as between the corrected predicted endpoint $\widehat{G}$ and the true endpoint $G$. The conditional variational autoencoder used for endpoint prediction is trained using Kullback-Leibler (KL) divergence as its loss function. The specific loss functions are defined in equations (29) and (30):
@@ III-D Loss Function
IV EXPERIMENTS
IV-A Dataset and Data Preprocessing
We trained and evaluated our model on two publicly available, large-scale, real-world highway datasets: NGSIM[37] [38] and HighD[39].
(1) NGIMS dataset: The NGSIM dataset was collected through a project initiated by the U.S. Federal Highway Administration. It contains real highway vehicle trajectory data from US101 and I80, with each sub-dataset capturing 45 minutes of trajectory data in mild, moderate, and congested traffic conditions. The dataset records vehicle position, speed, acceleration, and type at a sampling frequency of 10 Hz.
(2) HighD dataset: The HighD dataset was recorded using drones equipped with cameras over six highways in Germany, providing an aerial perspective. This dataset benefits from advanced machine learning algorithms, ensuring high localization accuracy. It includes data from over 110000 vehicles, covering a total mileage of 44500 kilometers, with information on vehicle position, speed, and acceleration sampled at 25 Hz.
Given the different sampling frequencies and data characteristics of the two datasets, we performed data preprocessing to standardize them. For our experiments, we used a 3-second historical trajectory to predict a 5-second future trajectory, creating 8-second scene segments during preprocessing. To reduce data volume, the sampling frequency was downsampled to 5 Hz, yielding 40 time steps per scenario, with 15 steps for history (${T_{h}}=15$) and 25 steps for prediction (${T_{f}}=25$). The datasets were divided into training, validation, and testing sets in a 7:1:2 ratio.
IV-B Experimental Setup and Evaluation
@@ IV-C Model Comparison
Figure 4: RSME Bar Chart of Each Model on NGSIM Dataset.
Figure 5: RSME Bar Chart of Each Model on HighD Dataset.
TABLE II: RSME Comparison of Each Model on NGSIM Dataset
@@
TABLE III: RSME Comparison of Each Model on HighD Dataset
@@
Tables II and III present the RMSE values for different models on the NGSIM and HighD datasets, respectively. A lower RMSE indicates smaller prediction errors. Bold text highlights the method with the best performance among all compared models. As shown in Tables II and III, our proposed EPN model achieved the best predictive performance on both the NGSIM and HighD datasets.
When compared to models such as L-LSTM, CS-LSTM, S-GAN, and WSiP, which only consider historical trajectory information, the prediction accuracy of EPN is significantly improved. This demonstrates that incorporating the vehicle’s driving status information and self-planning plays a crucial role in trajectory prediction. Furthermore, compared to the PiP model, which also accounts for the impact of ego vehicle planning on trajectory prediction, EPN exhibits a smaller prediction error, highlighting the benefits of our enhanced endpoint prediction module in improving model performance.In comparison with the S-TF model, which exhibits the smallest prediction error among the competing methods, EPN reduced the average RMSE by 34.9% on the NGSIM dataset and 64.6% on the HighD dataset. Figures 4 and 5 provide bar charts that visually compare the RMSE evaluation metrics across various models for both the NGSIM and HighD datasets. These results clearly demonstrate that our proposed EPN model outperforms the other models by a significant margin.
@@
Figure 6: ADE/FDE Bar Chart of Each Model on NGSIM Dataset.
Figure 7: ADE/FDE Bar Chart of Each Model on HighD Dataset.
TABLE IV: ADE/FDE Comparison of Each Model on NGSIM Dataset and HighD Dataset
@@
Table IV compares the ADE and FDE of EPN with the multimodal trajectory prediction methods CS-LSTM (M) and PiP (M). Figures 6 and 7 present the ADE and FDE evaluation metrics in the form of bar charts for the NGSIM and HighD datasets.
@@ IV-D Ablation Experiment
(5) EPN (R): Adds a endpoint prediction module to the PCS-LSTM model, while removing the endpoint correction mechanism from the full model.
@@
TABLE V: RSME/ADE/FDE Comparison of Ablation Experiments on NGSIM Dataset
@@
Table V presents the comparison results of RMSE, ADE, and FDE from the ablation experiments on the NGSIM dataset. Figure 8 provides a line graph that visually illustrates the impact of each module on the model’s predictive performance. Incorporating vehicle driving status information, such as speed and acceleration, improves trajectory prediction accuracy. Compared to the benchmark model, PCS-LSTM, the addition of speed as a feature in PCS-LSTM (V) results in average reductions of 6.6%, 11.7%, and 7.7% in RMSE, ADE, and FDE, respectively. Further including both speed and acceleration in PCS-LSTM (V+A) leads to average reductions of 10.6%, 16.1%, and 11.4% in these metrics. When comparing PCS-LSTM (V+A+M) to EPN (R), the latter shows an average reduction of 14.1%, 18.4%, and 21.5% in RMSE, ADE, and FDE, respectively. This highlights the significant impact of the endpoint prediction module on improving the performance of multimodal trajectory prediction. Finally, by comparing EPN (R) with the full EPN model, we observe that the addition of the endpoint correction mechanism contributes to further improvements in the model’s predictive performance.
(a) RSME line chart of ablation experiment in NGSIM dataset.
(b) ADE line chart of ablation experiment in NGSIM dataset.
(c) FDE line chart of ablation experiment in NGSIM dataset
Figure 8: Line graph of ablation experiment under NGSIM dataset.
@@
IV-E Visualization
In addition to the quantitative analysis of the model’s trajectory prediction performance, we also conducted a qualitative analysis. Figure 9 presents the multimodal trajectory prediction results of EPN in left turn, straight, and right turn driving scenarios for the target vehicle, along with a comparison of the optimal predictions between the EPN and PiP methods. Different colored curves are used to represent various trajectories: the gray curve shows the historical trajectories of all vehicles, the red curve represents the ego vehicle’s future planned trajectory, the green curve represents the true future trajectory of the target vehicle, the purple curve shows the predicted trajectory of the PiP method, and the blue curve indicates the predicted trajectory of EPN.
From the visualizations, it is clear that the multimodal trajectory predictions of EPN closely align with the true future trajectory of the target vehicle, demonstrating strong prediction performance. While the predictions for straight driving are similar between EPN and PiP, in left and right turn scenarios, the EPN predictions more closely match the true trajectory than those of PiP. This is because, in turning scenarios, predicting the vehicle’s possible endpoint directly—rather than relying on ambiguous driving intentions—enables a more accurate estimate of the vehicle’s movement, resulting in predictions that better match the true trajectory.
V CONCLUSION
In this paper, we propose a trajectory prediction model based on ego vehicle planning. This model uses the historical trajectory, vehicle speed, acceleration, and planned future trajectory of the ego vehicle as inputs, and outputs a multimodal prediction of the target vehicle’s future trajectory. By incorporating the planned trajectory of the ego vehicle, the interaction between the ego vehicle’s plan and the predicted trajectory of the target vehicle can be simulated, thereby more realistically replicating the driving scenario and improving the accuracy of trajectory prediction for the target vehicle. We introduce a target’s endpoint prediction module based on a CVAE to address the issues of intention ambiguity and large prediction errors typically found in trajectory prediction methods based on driving intentions. This module first predicts multiple potential endpoints for the target vehicle, and then refines the accuracy of these predictions through a endpoint correction mechanism. The module subsequently generates a complete multimodal trajectory based on the predicted endpoints. Experimental results demonstrate that our method achieves superior trajectory prediction accuracy compared to methods based solely on driving intentions.

### v2

@@ IV-C LSTM Decoder
In the LSTM decoder module, the endpoint encoder is used to encode the corrected endpoint $\widehat{G}$. The resulting encoded feature vector $\widehat{end}_{refine}$ is then concatenated with the environmental feature vector $enc$ and fed into the LSTM decoder. This process generates complete trajectories for each corrected endpoint $\widehat{G}$, resulting in multimodal trajectory prediction outcomes, the calculation process is shown in formulas (25) and (26):
$\widehat{end}_{refine}={E_{end}}(\widehat{G})$ (25)
$\widehat{Y}=LSTM(concat(enc,\widehat{end}_{refine}))$ (26)
@@
Due to significant variations in the absolute coordinate values of vehicles across different scenarios, our aim is to mitigate the adverse effects of these absolute values on the performance of trajectory prediction models. To achieve this, we predict the relative displacement $\left({\delta x_{t+\tau}^{i},\delta y_{t+\tau}^{i}}\right)$ of the target vehicle over the next ${T_{f}}$ time steps, where $\tau\in\left({1,2,...,{T_{f}}}\right)$. By predicting the relative displacement, we can improve the accuracy of the predictions. Once the relative displacement is predicted, the absolute coordinates of the predicted trajectory can be calculated using the formulas (27) and (28):
@@
IV-D Loss Function
We use the trajectory prediction error ${L_{pred}}$ and the CVAE error ${L_{cvae}}$ as loss functions. The trajectory prediction error ${L_{pred}}$ is calculated as the mean squared error (MSE) between the predicted trajectory $\widehat{Y}$ and the true trajectory $Y$, as well as between the corrected predicted endpoint $\widehat{G}$ and the true endpoint $G$. The CVAE used for endpoint prediction is trained using Kullback-Leibler (KL) divergence as its loss function. The specific loss functions are defined in equations (29) and (30):
@@ IV-D Loss Function
V EXPERIMENTS
V-A Dataset and Data Preprocessing
We trained and evaluated our model on two publicly available highway datasets: NGSIM [29], [30] and HighD [31].
(1) NGIMS dataset: The NGSIM dataset was collected through a project initiated by the U.S. Federal Highway Administration. The dataset records vehicle position, speed, acceleration, and type at a sampling frequency of 10 Hz.
(2) HighD dataset: The HighD dataset was recorded using drones with cameras over six highways in Germany, offering an aerial perspective. It includes vehicle position, speed, and acceleration sampled at 25 Hz.
Given the different sampling frequencies and data characteristics of the two datasets, we performed data pre-processing to standardize them. For our experiments, we used a 3-second historical trajectory to predict a 5-second future trajectory, creating 8-second scene segments during pre-processing. To reduce data volume, the sampling frequency was downsampled to 5 Hz, yielding 40 time steps per scenario, with 15 steps for history (${T_{h}}=15$) and 25 steps for prediction (${T_{f}}=25$). The datasets were divided into training, validation, and testing sets in a 7:1:2 ratio.
V-B Experimental Setup and Evaluation
@@ V-C Model Comparison
TABLE I: RSME Comparison of Each Model on NGSIM Dataset
@@
TABLE II: RSME Comparison of Each Model on HighD Dataset
@@
Tables I and II present the RMSE values for different models on the NGSIM and HighD datasets, respectively. A lower RMSE indicates smaller prediction errors. Bold text highlights the method with the best performance among all compared models. As shown in Tables I and II, our proposed EPN model achieved the best predictive performance on the NGSIM and HighD datasets.
When compared to models such as L-LSTM, CS-LSTM, S-GAN, and WSiP, which only consider historical trajectory information, the prediction accuracy of EPN is significantly improved. This demonstrates that incorporating the vehicle’s driving status information and ego vehicle planning plays a crucial role in trajectory prediction. Furthermore, compared to the PiP model, which also accounts for the impact of ego vehicle planning on trajectory prediction, EPN exhibits a smaller prediction error, highlighting the benefits of our enhanced endpoint prediction module in improving model performance. In comparison with the S-TF model, which exhibits the smallest prediction error among the competing methods, EPN reduced the average RMSE by 34.9% on the NGSIM dataset and 64.6% on the HighD dataset. These results clearly demonstrate that our proposed EPN model outperforms the other models by a significant margin.
@@
TABLE III: ADE/FDE Comparison of Each Model on NGSIM Dataset and HighD Dataset
@@
Table III compares the ADE and FDE of EPN with the multimodal trajectory prediction methods CS-LSTM (M) and PiP (M).
@@ V-D Ablation Experiment
(5) EPN (R): Adds an endpoint prediction module to the PCS-LSTM model, while removing the endpoint correction mechanism from the full model.
@@
TABLE IV: RSME/ADE/FDE Comparison of Ablation Experiments on NGSIM Dataset
@@
Table IV presents the comparison results of RMSE, ADE, and FDE from the ablation experiments on the NGSIM dataset. Incorporating vehicle driving status information, such as speed and acceleration, improves trajectory prediction accuracy. Compared to the benchmark model, PCS-LSTM, the addition of speed as a feature in PCS-LSTM (V) results in average reductions of 6.6%, 11.7%, and 7.7% in RMSE, ADE, and FDE, respectively. Further including both speed and acceleration in PCS-LSTM (V+A) leads to average reductions of 10.6%, 16.1%, and 11.4% in these metrics. When comparing PCS-LSTM (V+A+M) to EPN (R), the latter shows an average reduction of 14.1%, 18.4%, and 21.5% in RMSE, ADE, and FDE, respectively. This highlights the significant impact of the endpoint prediction module on improving the performance of multimodal trajectory prediction. Finally, by comparing EPN (R) with the full EPN model, we observe that the addition of the endpoint correction mechanism contributes to further improvements in the model’s predictive performance.
@@
V-E Visualization
We visualize the trajectory data in different scenarios to qualitatively analyze the performance of the model. Figure 3 presents the multimodal trajectory prediction results of the EPN in left turn, straight, and right turn driving scenarios for the target vehicle, along with a comparison of the optimal predictions between the EPN and PiP methods. Different colored curves are used to represent various trajectories.
From the visualizations, it is clear that the multimodal trajectory predictions of EPN closely align with the true future trajectory of the target vehicle, demonstrating strong prediction performance. While the predictions for straight driving are similar between EPN and PiP, in left and right turn scenarios, the EPN predictions more closely match the true trajectory than those of PiP. This is because, in turning scenarios, predicting the vehicle’s possible endpoint directly, rather than relying on ambiguous driving intentions, enables a more accurate estimate of the vehicle’s movement, resulting in predictions that better match the true trajectory.
VI CONCLUSION
In this paper, we propose a trajectory prediction model based on ego vehicle planning. This model uses the historical trajectory, vehicle speed, acceleration, and planned future trajectory of the ego vehicle as inputs, and outputs a multimodal prediction of the target vehicle’s future trajectory. By incorporating the planned trajectory of the ego vehicle, the interaction between the ego vehicle’s plan and the predicted trajectory of the target vehicle can be simulated, thereby more realistically replicating the driving scenario and improving the accuracy of trajectory prediction for the target vehicle. We introduce a target’s endpoint prediction module based on a CVAE to address the issues of intention ambiguity and large prediction errors typically found in trajectory prediction methods based on driving intentions. This module first predicts multiple potential endpoints for the target vehicle, and then refines the accuracy of these predictions through an endpoint correction mechanism. The module subsequently generates a complete multimodal trajectory based on the predicted endpoints. Experimental results demonstrate that our method achieves superior trajectory prediction accuracy compared to methods based solely on driving intentions.